12-3(x=1)=+17
To solve I first multiply 3 timesx then move x from right to left then add 12 and 17, I am geeting confused. Please help me solve.
Thank you

Answer by Earlsdon(6294)

Solve for x:
12-3(x+1) = 17
Did I get this right?
Apply the "distributive property"
12-3x-3 = 17 Simplify this.
9-3x = 17 Subtract 9 from both sides.
-3x = 8 Finally, divide both sides by -3.
x = -8/3
Check: Substitute x = -8/3
12-3(-8/3+1) = 17
12-(-8+3) = 17
12-(-5) = 17
12+5 = 17
17 = 17 It checks!